Find the derivative of y = ln 3x 5


1
Expert's answer
2021-06-25T07:13:08-0400
"y=\ln(3x^5)=\ln(3)+5\ln(x)"

"y'(x)=(\ln(3x^5))'"

"=(\ln(3)+5\ln(x))'=\dfrac{5}{x}, x>0"